OpenStudy (aye_my_peeps):

solve for x answer choices: a. -3 b. 1 c.-14 d. -10

OpenStudy (k12girlohva):

Isn there a number sentence or something

OpenStudy (kendricklamar2014):

@AYE_MY_PEEPS where is the math problem?

OpenStudy (aye_my_peeps):

\[\frac{ X+2 }{ 3 }= -4\]

OpenStudy (kendricklamar2014):

Simplify: \[\frac{ 1 }{ 3 }x + \frac{ 2 }{ 3 } = -4\] Step 2: Subtract 2/3 from both sides: \[\frac{ 1 }{ 3 }x + \frac{ 2 }{ 3 } - \frac{ 2 }{ 3 } = -4 - \frac{ 2 }{ 3 }\] \[= \frac{ 1 }{ 3 }x = \frac{ -14 }{ 3 }\]

OpenStudy (kendricklamar2014):

Step 3: Multiply both sides by 3: \[3 \times \frac{ 1 }{ 3 }x = 3 \times \frac{ -14 }{ 3 }\] \[x =\]

OpenStudy (aye_my_peeps):

SO -14??

OpenStudy (aye_my_peeps):

C??

OpenStudy (kendricklamar2014):

Yes, thats the answer